# From Exploration to Mastery: Enabling LLMs to Master Tools via Self-Driven Interactions

### v1

@@ 4 Related Work
Tool Learning. Recent studies have highlighted the potential of LLMs to utilize external tools in addressing complex problems (Qu et al., 2024b; Wang et al., 2024d). With the aid of external tools, LLMs can obtain up-to-date information (Nakano et al., 2021; Gou et al., 2024a; b), enhance domain-specific knowledge (M. Bran et al., 2024; Zhang et al., 2024a), process multi-modal information (Surís et al., 2023; Gao et al., 2024c), and more. Existing tool learning approaches can be categorized into two types: tuning-based and tuning-free methods (Gao et al., 2024b). Tuning-based methods enhance the tool-using capabilities of LLMs by fine-tuning them on tool-related datasets (Patil et al., 2023; Hao et al., 2024; Yang et al., 2024; Qin et al., 2024; Liu et al., 2024). However, this approach is only applicable to open-source models and requires substantial computational resources. In contrast, tuning-free methods provide LLMs with tool documentation and a few demonstrations (Wei et al., 2022; Hsieh et al., 2023; Paranjape et al., 2023; Du et al., 2024; Shi et al., 2024), relying on the in-context learning ability of LLMs to understand how to use tools. This approach requires no additional training and allows for the plug-and-play integration of external tools. However, this method necessitates high-quality tool documentation that is aligned with the comprehension of LLMs (Yuan et al., 2024). In this paper, we propose a method to align with LLMs understanding and improve the quality of tool documentation to enhance the tool-using capabilities of LLMs.
@@ 5 Conclusion
Appendix A More Experiments
We also conduct analysis experiments on three datesets using other LLMs. As illustrated in Table 5, both GPT-4o-mini and LLama-3-70B display trends that are consistent with those observed in GPT-4o. The absence of our proposed diversity-promoting exploration strategy and tool-adaptive termination mechanism results in a decline in performance, thereby underscoring the importance of our design innovations. An intuitive explanation posits that, in the absence of the diversity-promoting exploration strategy, the examples generated during each round of exploration may exhibit significant similarity. This similarity could render multiple iterations indistinguishable from a single iteration, thereby undermining the advantages of repeated exploration. Likewise, the absence of the tool-adaptive termination mechanism may result in excessive iterations, which could produce redundant information and contribute to overfitting.
@@ Appendix A More Experiments
Appendix B Detailed Prompts
@@
Appendix C Case Study
C.1 ToolBench
@@ C.2 RestBench
Appendix D Revision Trajectory
@@
Through an analysis of the revision trajectory, it becomes evident that during the first round, the experience of encountering an invalid ID is documented. The analyzer promptly identifies the issue and provids suggestions for improvement. Subsequently, the rewriter utilizes this feedback to make appropriate updates to the tool documentation, recommending that future exploration should concentrate on retrieving results associated with a valid ID. By the third iteration, the exploration successfully yields results with a valid ID, leading to further updates in the documentation. This process underscores the necessity of conducting multiple iterations. Furthermore, in the absence of a diversity-promoting exploration strategy designed to facilitate diverse exploration, the system may repeatedly encounter similar scenarios, thereby hindering the discovery of valid IDs and obstructing effective updates. This situation emphasizes the critical importance of implementing a diversity-promoting exploration strategy.

### v2

@@ 4 Related Work
Tool Learning. Recent studies have highlighted the potential of LLMs to utilize external tools in addressing complex problems (Qu et al., 2024b; Wang et al., 2024d). With the aid of external tools, LLMs can obtain up-to-date information (Nakano et al., 2021; Gou et al., 2024a; b), enhance domain-specific knowledge (M. Bran et al., 2024; Zhang et al., 2024a), process multi-modal information (Surís et al., 2023; Gao et al., 2024c), and more. Existing tool learning approaches can be categorized into two types: tuning-based and tuning-free methods (Gao et al., 2024b). Tuning-based methods enhance the tool-using capabilities of LLMs by fine-tuning them on tool-related datasets (Patil et al., 2023; Hao et al., 2024; Yang et al., 2024; Qin et al., 2024; Liu et al., 2024). However, this approach is only applicable to open-source models and requires substantial computational resources. In contrast, tuning-free methods provide LLMs with tool documentation and a few demonstrations (Wei et al., 2022; Hsieh et al., 2023; Paranjape et al., 2023; Du et al., 2024; Shi et al., 2024), relying on the in-context learning ability to understand how to use tools. This approach requires no additional training and allows for the plug-and-play integration of external tools. However, this method necessitates high-quality tool documentation that is aligned with the comprehension of LLMs (Yuan et al., 2024; Chen et al., 2024). In this paper, we propose a method to align with LLMs understanding and improve the quality of tool documentation to enhance the tool-using capabilities of LLMs.
@@ 5 Conclusion
Acknowledgements
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (62472426), PCC@RUC, fund for building world-class universities (disciplines) of Renmin University of China. Work partially done at Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education.
@@ Acknowledgements
Appendix A Details of Human Evaluation
We recruit three doctoral students familiar with the domain to evaluate the raw tool documentation and its revised versions generated by DRAFT and pay accordingly. They are selected based on their familiarity with the domain and their capability to assess the quality of tool documentation accurately. They are asked to evaluate the documentation through pair-wise comparison based on the following three criteria: 1. Completeness: Which documentation more thoroughly describes the tool’s potential uses and key information, such as its usage and parameter descriptions, without missing critical details. 2. Conciseness: Which documentation is clearer and more concise, avoiding irrelevant or redundant information. 3. Accuracy: Which documentation describes the tool functionality more accurately and is free from incorrect information. The Fleiss’ Kappa statistics for completeness, conciseness, and accuracy are 0.76, 0.71, and 0.72 on RestBench, while 0.77, 0.69, and 0.71 on ToolBench, indicating a high agreement among the three annotators.
Appendix B More Experiments
We also conduct analysis experiments on three datasets using other LLMs. As illustrated in Table 5, both GPT-4o-mini and LLama-3-70B display trends that are consistent with those observed in GPT-4o. The absence of our proposed diversity-promoting exploration strategy and tool-adaptive termination mechanism results in a decline in performance, thereby underscoring the importance of our design innovations. An intuitive explanation posits that, in the absence of the diversity-promoting exploration strategy, the examples generated during each round of exploration may exhibit significant similarity. This similarity could render multiple iterations indistinguishable from a single iteration, thereby undermining the advantages of repeated exploration. Likewise, the absence of the tool-adaptive termination mechanism may result in excessive iterations, which could produce redundant information and contribute to overfitting.
@@ Appendix B More Experiments
Appendix C Detailed Prompts
@@
Appendix D Case Study
D.1 ToolBench
@@ D.2 RestBench
Appendix E Revision Trajectory
@@
Through an analysis of the revision trajectory, it becomes evident that during the first round, the experience of encountering an invalid ID is documented. The analyzer promptly identifies the issue and provides suggestions for improvement. Subsequently, the rewriter utilizes this feedback to make appropriate updates to the tool documentation, recommending that future exploration should concentrate on retrieving results associated with a valid ID. By the third iteration, the exploration successfully yields results with a valid ID, leading to further updates in the documentation. This process underscores the necessity of conducting multiple iterations. Furthermore, in the absence of a diversity-promoting exploration strategy designed to facilitate diverse exploration, the system may repeatedly encounter similar scenarios, thereby hindering the discovery of valid IDs and obstructing effective updates. This situation emphasizes the critical importance of implementing a diversity-promoting exploration strategy.